OpenStudy (anonymous):

Algebra help please! I do not get this..... Simplify the expression.. \[27^\frac{ 1 }{ 3 }\] Use the rules for exponents simplify the expression. \[b\frac{ m }{ n }=n \sqrt{b}^m=(n \sqrt{b})^m\]

terenzreignz (terenzreignz):

As if to support L66 ^_^ \[\huge a^{\frac{\color{red}m}{\color{green}n}}=\sqrt[\color{green}n]{a^{\color{red}m}}\]

OpenStudy (loser66):

ok, so let TJ help

terenzreignz (terenzreignz):

Look closely at what happens to the numerator-exponent and denominator-exponent.

terenzreignz (terenzreignz):

Oh no, I was only (technically) cheering for you L66 :) Please continue XD

OpenStudy (loser66):

TJ I know you, you have a good way to explain, please, please, help him

terenzreignz (terenzreignz):

Fine :3 @Dakotafox79 I want you to factor 27, completely, into primes... Can you do that?

OpenStudy (anonymous):

Erm, no. Sorrry for slow response laggy school pc....

terenzreignz (terenzreignz):

You can't factor 27? :| Very well, let's not do that XD Use this rule... \[\huge a^{\frac{\color{red}m}{\color{green}n}}=\sqrt[\color{green}n]{a^{\color{red}m}}\] And in this instance... \[\Huge 27^{\frac{\color{red}1}{\color{green}3}}=\color{blue}? \]

terenzreignz (terenzreignz):

What do you get? Use the drawing tool, if you don't know how to typeset equations yet ^_^

terenzreignz (terenzreignz):

Hey, @Dakotafox79 Stuck?

OpenStudy (anonymous):

\[\sqrt[3]{27}^1\] Is that it ? Lol... God I suck at math. :/

terenzreignz (terenzreignz):

That is correct :) \[\Huge 27^{\frac{\color{red}1}{\color{green}3}}=\sqrt[\color{green}3]{27^\color{red}1}=\sqrt[3]{27}\] L66 has taught you well... Now... what's the cube root of 27? ^_^

OpenStudy (anonymous):

3, right?

terenzreignz (terenzreignz):

Brilliant :)

terenzreignz (terenzreignz):

Now that wasn't so hard, was it?

OpenStudy (anonymous):

Yay, thank you so much :D. And now that I look at it like this, not really that hard :) thanks so much!!!

terenzreignz (terenzreignz):

Of course. Might help you to remember that... raising to an exponent 1/2 is the same as getting the square root... raising to an exponent 1/3 is the same as getting cube root raising to an exponent 1/4 is the same as getting fourth root... etc It'll save you time ^_^

OpenStudy (anonymous):

Okay. Thanks again :)